# Differentially Private Random Feature Model

### v1

@@ 3.3 Generalization Error
with probability at least $21-\delta/$ 2. It is obvious to see that this upper bound is worse than the bound in Lemma 10 in terms of privacy budget $\epsilon_{p}$ and number of features $N$. Our experimental results also support this observation, see comparison in Section 4.
3.4 Removing disparate impact
In this section, we aim to show that random feature model has the potential to remove disparate impact. It has been observed that the unfairness, such as disparate impact, exists in the DP mechanism [28, 2, 3]. The disparate impact refers to the phenomenon that the performance of a differentially private model for underrepresented classes and subgroups tends to be worse than that of the original, non-private model. We adopt the notion of excessive risk gap as the fairness metric, which has been used in fairness analysis, see [27]. To define the excessive risk gap, we first define excessive risk, which is a widely adopted metric to measure utility in private learning [38, 39]. Consider the regression problem (it can be generalized to any supervised learning problem) with mean-squared loss, we define the risk function $L(\theta,D)$ for any given dataset $D$ as
@@ 3.4 Removing disparate impact
Then, we define the excessive risk, which is the difference between the private and non-private risk functions, as
where the expectation is defined over the randomness of the private mechanism and $\hat{\theta}$ denotes the private model parameter while $\theta^{\sharp}=\mathop{\mathrm{argmin}}_{\theta}L(\theta,D)$. In this paper, we use $R(\theta)$ and $R_{a}(\theta)$ to denote, respectively, the population-level $R(\theta,D)$ excessive risk and the group level $R(\theta,D_{a})$ excessive risk for sensitive group $a$. Then, we defined the excessive risk gap for each group $a$ as
$\xi_{a}=|R(\theta)-R_{a}(\theta)|.$
Fairness is achieved when $\xi_{a}=0$ for all groups $a$. Unfairness occurs when different groups have different excessive risk gaps. Moreover, a larger excessive risk gap for a certain group implies that private estimator leads to a larger utility loss for this group. Following the analysis in [27], for a fixed privacy budget $\epsilon$, the excessive risk gap for linear models mainly depends on the trace of Hessian matrix $\mathbf{H}^{a}=\nabla_{\theta^{\sharp}}^{2}L(\theta^{\sharp},D_{a})$ of risk function, at the optimal parameter vector $\theta^{\sharp}$. Suppose that the Gaussian mechanism is applied to produce $(\epsilon_{p},\delta_{p})$-differentially private estimator $\hat{\theta}$, i.e adding noise drawn from normal distribution $\mathcal{N}(0,\sigma^{2})$ to each entry of the optimal model parameter $\theta^{\sharp}$. Theorem 1 in [27] indicates that the excessive risk gap $\xi_{a}$ is approximated by
@@
Moreover, the trace of the Hessian matrix $\mathbf{H}^{a}$ depends solely on the input norms of the elements in $D_{a}$ for each group $a$ since $\mathrm{\mathop{Tr}}(\mathbf{H}^{a})=\mathbb{E}_{\mathbf{x}\in D_{a}}\|\mathbf{x}\|^{2}$, which highlights the relation between fairness and the average input norms of different group elements. When these norms are substantially different from one another they will impact their respective excessive risks differently. Based on this observation, Corollary 2 in [27] suggests that better fairness may be achieved by normalizing the input values for each group independently in the regression setting.
The random feature method can be viewed as a normalization approach using random projections, and therefore it helps to achieve better fairness. Since we generate a set of random features $\{\boldsymbol{\omega}_{k}\}_{k\in[N]}\subset\mathbb{R}^{d}$ and create new features by mapping $\mathbf{x}$ to
@@ 4 Numerical Experiments
In this section, we present numerical experiments on the DP random feature model in Algorithm 1. We divide this section into two parts: i) we compare our proposed method with the regression baselines, and ii) we show that the DP random feature model has the potential to reduce disparate impact. The code is available on Github repository: https://github.com/liaochunyang/DPRFM
@@ 4.1 Generalization performance
Our goal of this section is to compare the generalization performance of our proposed method with the regression baselines in [20] and [25]. The generalization performance is evaluated on a hold-out test set of size $N_{test}$ and the test error is defined by
$TestError=\frac{1}{N_{test}}\sum_{i=1}^{N_{test}}(f(\mathbf{x}_{i})-\hat{f}(\mathbf{x}_{i}))^{2}.$
We test our algorithm on both synthetic and real data. Throughout the experiments, we consider the Gaussian random feature models where random feature $\boldsymbol{\omega}\in\mathbb{R}^{d}$ is drawn from a multivariate normal distribution $\mathcal{N}(0,\sigma^{2}\mathbf{I}_{d})$. The choice of variance $\sigma^{2}$ varies among each example and will be stated clearly.
Differences between regression baselines. In our numerical experiments, we mainly compare our proposed method with the regression baselines in [20] and [25]. The difference between our method and that in [20] is that different random noises are added to produce private estimators. Our method adds Gaussian random noise, while the existing method in [20] adds random noise generated from a probability distribution with density function $\propto\exp(-\xi\|b\|_{2})$. Our theory suggests that the private model with Gaussian random noise has better generalization performance. The numerical experiments verify our theory empirically. In [25], the authors studied a private random feature model via SGD. However, the non-private solution solved by using SGD may not interpolate the training data due to the restrictions on the step size and total number of iterations. Our method focuses on the interpolation regime and then produces a private estimator. In our numerical experiments, the methods in [20] and [25] will be referred to as Gamma and SGD, respectively. (The former is called Gamma since the $\ell_{2}$ norm of additive random noise follows a Gamma distribution.)
The setting of SGD follows the set-up from [25]. Specifically, the initialization will be the zero vector, the batch size will be 1, the step size(learning rate) is selected to be $1/m$, and the number of iterations is $m$, where $m$ is the training samples size.
@@
Real data. For our experiments on real data, we use the medical cost dataset [42] and the wine quality dataset [43]. The medical cost dataset looks at an individual medical cost prediction task. Each individual has three numeric {age, BMI, #children} and three categorical features {sex, smoker, region}. The dataset also has a real-valued medical charges column that we use as label. We use min-max scaling to normalize the numeric features as well as the label to be within the range $[0,1]$. For any categorical features, we use standard one-hot-encoding. The wine quality dataset is used to model wine quality based on physicochemical tests, which contains 11 numeric features and 1 label column named quality. There are no categorical features in this dataset. Similarly, we use min-max scaling to normalize the numeric features as well as the label.
@@
We also compare our model with the regression baseline proposed in [25] in terms of test error. To implement SGD, we take the batch size to be 1, number of iterations to be $m$, and step sizes $\eta=\{1/m,1/(2m),1/(8m)\}$ as suggested in [25], where $m$ is the sample size. The min-norm interpolation problem is solved by using the randomized Kaczmarz algorithm, which is a special case of SGD. The number of iterations is selected to be $m$ as well. The results are shown in Figure 5 (Medical cost dataset) and Figure 6 (Wine quality dataset). We first compare the performances of non-private random feature models produced by solving the min-norm interpolation problem and by using SGD. Our results indicate that the min-norm solution can generalize well and achieve better test performance than models trained by SGD. We also observe that SGD (with other step sizes) does not guarantee the convergence to the min-norm solution. On the comparisons of private random feature models, our experiments show that our proposed method has better test performance than the regression baseline.
@@
Figure 5: Medical cost dataset: we compare test errors of non-private random feature models (on the left) and corresponding private random feature models (on the right). Different step sizes are considered when using SGD. Gaussian random features are generated from normal distribution with mean 0 and variance $\sigma^{2}=2$. The privacy budget $\epsilon_{p}$ is 0.5 and the privacy parameter $\delta_{p}=10^{-5}$.
@@
Figure 6: Wine quality dataset: we compare test errors of non-private random feature models (on the left) and corresponding private random feature models (on the right). Different step sizes are considered when using SGD. Gaussian random features are generated from normal distribution with mean 0 and variance $\sigma^{2}=2$. The privacy budget $\epsilon_{p}$ is 0.5 and the privacy parameter $\delta_{p}=10^{-5}$.
@@
4.2 Disparate impact
The goal of this section is to show that the DP random feature model does not seem to suffer disparate impact empirically. We still use the medical cost and wine quality datasets for this section. We compare our private random feature model with the private regularized linear regression model proposed in [20, 37].
Data processing. We use standard one-hot-encoding for categorical features and normalize labels $y$ by dividing its $\ell_{2}$ norm. For the random feature model, we do not normalize any numerical features since we want to make sure that the changes in the excessive risk gap do not depend on data normalization. For the regularized linear regression model, we normalize the input values for each group independently. In the wine quality experiment, group 1 and 2 represent red wine and white wine, respectively. We randomly sample 1000 data points from each group for training and 500 samples from each group for testing. In the medical cost experiment, we first use categorical feature sex to create two different groups, i.e. group 1 means male and group 2 means female. In the second experiment, we divide the dataset into two groups according to the categorical feature smoker, where group 1 means smoker and group 2 represents non-smoker. In those examples, 90% of the data are randomly sampled from each group for training and the remaining 10% of data are for testing.
@@ 4.2 Disparate impact
(a) Medical cost - Example 1
(b) Medical cost - Example 2
Figure 7: Training excessive risk gaps for each group versus privacy budget $\epsilon_{p}$ for private random feature model and regularized linear regression model. We consider medical cost dataset in this experiment. On the left: ”sex” is selected as sensitive feature. On the right: ”smoker” is the sensitive feature.
Figure 8: Training excessive risk gaps for each group versus privacy budget $\epsilon_{p}$ for private random feature model and regularized linear regression model. The wine quality dataset is considered here.
Figure 7 and 8 show the relations between the excessive risk gaps for each group and the machine learning models while varying the privacy budget $\epsilon_{p}$. Each data point represents the average of 10 runs of a train-test split. In all experiments, the excessive risk gaps for each group of the random feature model are smaller than excessive risk gaps of the regularized linear regression model (provided the same privacy budget). This means that the random feature model does not lead to utility loss. Moreover, the excessive risk gaps across different groups are similar for the random feature model, especially when the privacy budget $\epsilon_{p}$ is small, which indicates that better fairness may be achieved by using the private random feature model. In addition, we observe that the excessive risk gap decreases as the privacy budget $\epsilon_{p}$ increases. The same phenomenon has been observed in [27]. This is likely due to the fact that larger $\epsilon_{p}$ values require smaller $\sigma$ values, and thus, as shown in equation (8), the excessive risk gap decreases for a fixed Hessian trace. In summary, we believe that better fairness may be achieved by using the random feature model and it is thus superior to feature normalization.
@@ 5 Discussion
In this paper, we established theoretical guarantees for the differentially private random feature model. The random feature model is a widely used surrogate model for differentially private learning with kernels. Apart from the current literature where people focused on the regularization regime, we focused on the overparametrized regime and studied the min-norm interpolation estimator. Our setting is closer to the modern deep learning setting where the number of trainable parameters is much larger than the number of samples and explicit regularization is not added. Our privacy guarantee follows the standard Gaussian mechanism and the derivation of the generalization error bound relies on tools from random matrix theory. We also show that the proposed DP random feature model seems to not suffer disparate impact. Additionally, numerical experiments showed that our method has better generalization performance under the same privacy budget compared with the current differentially private random feature method, and has the potential to reduce disparate impact.
We conclude with some directions for future work. First, it would be interesting to study differentially private shallow neural networks through the random feature model as it is an approximation of a shallow neural network. Second, our results allow us to consider differentially private kernel interpolation models. Moreover, current literature focuses on shift-invariant kernels, such as the Gaussian kernel, because the random features for shift-invariant kernels are easy to construct. The generalization to an arbitrary kernel, e.g. polynomial kernels, deserves further investigation. Finally, we studied the disparate impact of the DP random feature model under the regression setting in this paper through the excessive risk gap. Other choices of fairness constraints for regression problems, such as statistical parity used in [6], should be studied as well. Understanding the fairness properties of the DP random feature model in the classification setting will also be interesting future work.
@@
We first give a bound of $\|f-f^{*}\|_{L^{2}(\mu)}$. As a consequence of Lemma 11 and 12, we can conclude that
@@
[…]
@@
$v(\mathbf{u}_{1},\dots,\mathbf{u}_{m})$ $:=\|f^{*}-f^{\sharp}\|^{2}_{L^{2}(\mu)}-\frac{1}{m}\sum_{j=1}^{m}\left|f^{*}(\mathbf{u}_{j})-f^{\sharp}(\mathbf{u}_{j})\right|^{2}$
@@
$\leq$ $\frac{2N}{m}\|\mathbf{c}^{*}-\mathbf{c}^{\sharp}\|_{2}^{2}:=\Delta_{v},$

### v2

@@ 3.3 Generalization Error
with probability at least $1-\delta/2$. It is obvious to see that this upper bound is worse than the bound in Lemma 13 in terms of privacy budget $\epsilon_{p}$ and number of features $N$. Our experimental results also support this observation, see comparison in Section 4.
3.4 Fairness Guarantee
In this section, we aim to show that random feature model has the potential to remove disparate impact via the notation of excessive risk gap. It has been observed that the unfairness, such as disparate impact, exists in the DP mechanism [2, 3, 31]. The disparate impact refers to the phenomenon that the performance of a differentially private model for underrepresented classes and subgroups tends to be worse than that of the original, non-private model. Consider the regression problem (it can be generalized to any supervised learning problem) with mean-squared loss, we can explicit write the risk function (1) as
@@ 3.4 Fairness Guarantee
Following the analysis in [28], for a fixed privacy budget $\epsilon$, the excessive risk gap for linear models mainly depends on the trace of Hessian matrix $\mathbf{H}^{a}=\nabla_{\theta^{\sharp}}^{2}L(\theta^{\sharp},D_{a})$ of risk function, at the optimal parameter vector $\theta^{\sharp}$. Suppose that the Gaussian mechanism is applied to produce $(\epsilon_{p},\delta_{p})$-differentially private estimator $\hat{\theta}$, i.e adding noise drawn from normal distribution $\mathcal{N}(0,\sigma^{2})$ to each entry of the optimal model parameter $\theta^{\sharp}$. Theorem 1 in [28] indicates that the excessive risk gap $\xi_{a}$ (see Definition 4) is approximated by
@@
Moreover, the trace of the Hessian matrix $\mathbf{H}^{a}$ depends solely on the input norms of the elements in $D_{a}$ for each group $a$ since $\mathrm{\mathop{Tr}}(\mathbf{H}^{a})=\mathbb{E}_{\mathbf{x}\in D_{a}}\|\mathbf{x}\|^{2}$, which highlights the relation between fairness and the average input norms of different group elements. When these norms are substantially different from one another they will impact their respective excessive risks differently. Based on this observation, Corollary 2 in [28] suggests that better fairness may be achieved by normalizing the input values for each group independently in the regression setting. The random feature method can be viewed as a normalization approach using random projections, and therefore it helps to achieve better fairness. Since we generate a set of random features $\{\boldsymbol{\omega}_{k}\}_{k\in[N]}\subset\mathbb{R}^{d}$ and create new features by mapping $\mathbf{x}$ to
@@ 4 Numerical Experiments
In this section, we present numerical experiments on the DP random features model in Algorithm 1. We divide this section into two parts: i) we compare our proposed method with the several regression baselines in the literature, and ii) we numerically verify that the DP random feature model has the potential to reduce disparate impact via the notations of excessive risk gap and statistical parity. The code is available on Github repository: https://github.com/liaochunyang/DPRFM
@@ 4.1 Generalization performance
In this section, we compare the generalization performance of our proposed method with the regression baselines in [22] and [29]. The generalization performance is evaluated on a hold-out test set $D_{test}$ of size $|D_{test}|$. We define the test error as
$TestError=\frac{1}{|D_{test}|}\sum_{(\mathbf{x}_{i},y_{i})\in D_{test}}(y_{i}-\hat{f}(\mathbf{x}_{i}))^{2},$
where $\hat{f}$ is an approximator. We test our algorithm on both synthetic and real data. Throughout the experiments, we consider the Gaussian random feature models where random feature $\boldsymbol{\omega}\in\mathbb{R}^{d}$ is drawn from a multivariate normal distribution $\mathcal{N}(0,\sigma^{2}\mathbf{I}_{d})$. The choice of variance $\sigma^{2}$ varies among each example and will be stated clearly.
Differences between regression baselines. In our numerical experiments, we mainly compare our proposed method with the regression baselines in [22] and [29]. The difference between our method and that in [22] is that different random noises are added to produce private estimators. Our method adds Gaussian random noise, while the existing method in [22] adds random noise generated from a probability distribution with density function $\propto\exp(-\xi\|b\|_{2})$. Our theory suggests that the private model with Gaussian random noise has better generalization performance. The numerical experiments verify our theory empirically. In [29], the authors studied a private random feature model via SGD. However, the non-private solution solved by using SGD may not interpolate the training data due to the restrictions on the learning rate and total number of iterations. Our method focuses on the interpolation regime and then produces a private estimator. In our numerical experiments, the methods in [22] and [29] will be referred to as Gamma and SGD, respectively. (The former is called Gamma since the $\ell_{2}$ norm of additive random noise follows a Gamma distribution.)
The setting of SGD follows the set-up from [29]. Specifically, the initialization will be the zero vector, the batch size will be 1, the learning rate is selected to be $1/m$, and the number of iterations is $m$, where $m$ is the training samples size.
@@
Real data. For our experiments on real data, we use the medical dataset [44] and the wine quality dataset [45]. The medical dataset looks at an individual medical cost prediction task. Each individual has three numeric {age, BMI, #children} and three categorical features {sex, smoker, region}. The dataset also has a real-valued medical charges column that we use as label. We use min-max scaling to normalize the numeric features as well as the label to be within the range $[0,1]$. For any categorical features, we use standard one-hot-encoding. The wine quality dataset is used to model wine quality based on physicochemical tests, which contains 11 numeric features and 1 label column named quality. There are no categorical features in this dataset. Similarly, we use min-max scaling to normalize the numeric features as well as the label.
@@
We also compare our model with the regression baseline proposed in [29] in terms of test error. To implement SGD, we take the batch size to be 1, number of iterations to be $m$, and learning rate $\lambda=\{1/m,1/(2m),1/(8m)\}$ as suggested in [29], where $m$ is the sample size. The min-norm interpolation problem is solved by using the randomized Kaczmarz algorithm, which is a special case of SGD. The number of iterations is selected to be $m$ as well. The results are shown in Figure 5 (Medical cost dataset) and Figure 6 (Wine quality dataset). We first compare the performances of non-private random feature models produced by solving the min-norm interpolation problem and by using SGD. Our results indicate that the min-norm solution can generalize well and achieve better test performance than models trained by SGD. We also observe that SGD (with other learning rates) does not guarantee the convergence to the min-norm solution. On the comparisons of private random feature models, our experiments show that our proposed method has better test performance than the regression baseline.
@@
Figure 5: Medical cost dataset: we compare test errors of non-private random feature models (on the left) and corresponding private random feature models (on the right). Differentlearning rates $\lambda$ are considered when using SGD. Gaussian random features are generated from normal distribution with mean 0 and variance $\sigma^{2}=2$. The privacy budget $\epsilon_{p}$ is 0.5 and the privacy parameter $\delta_{p}=10^{-5}$.
@@
Figure 6: Wine quality dataset: we compare test errors of non-private random feature models (on the left) and corresponding private random feature models (on the right). Different learning rates $\lambda$ are considered when using SGD. Gaussian random features are generated from normal distribution with mean 0 and variance $\sigma^{2}=2$. The privacy budget $\epsilon_{p}$ is 0.5 and the privacy parameter $\delta_{p}=10^{-5}$.
@@
4.2 Excessive Risk Gap
In this section, we numerically show that our proposed DP random features model achieves smaller excessive risk gap resulting in fairer results. We use the medical cost and wine quality datasets for this section and compare our proposed DP random features model with the private regularized linear regression model proposed in [22, 41].
Data processing. We use standard one-hot-encoding for categorical features and normalize labels $y$ by dividing its $\ell_{2}$ norm. For the random feature model, we do not normalize any numerical features since we want to make sure that the changes in the excessive risk gap do not depend on data normalization. For the regularized linear regression model, we normalize the input values for each group independently. In the wine quality experiment, group 1 and 2 represent red wine and white wine, respectively. We randomly sample 1000 data points from each group for training and 500 samples from each group for testing. In the medical cost experiment, we first use categorical feature sex to create two different groups, i.e. group 1 is male group and group 2 is female group. In the second experiment, we divide the dataset into two groups according to the categorical feature smoker, where group 1 is smoker group and group 2 is non-smoker group. In those examples, 90% of the data are randomly sampled from each group for training and the remaining 10% of data are for testing.
@@ 4.2 Excessive Risk Gap
(a) Sensitive feature ”sex”
(b) Sensitive feature ”smoker”
Figure 7: Medical cost dataset: Training excessive risk gaps for each group versus privacy budget $\epsilon_{p}$ for private random features model and regularized linear regression model.
Figure 8: Wine quality dataset: Training excessive risk gaps for each group versus privacy budget $\epsilon_{p}$ for private random features model and regularized linear regression model.
Figure 7 and 8 show the relations between the excessive risk gaps for each group and the machine learning models while varying the privacy budget $\epsilon_{p}$. Each data point represents the average of 10 runs of a train-test split. In all experiments, the excessive risk gaps for each group of the random feature model are smaller than excessive risk gaps of the regularized linear regression model (provided the same privacy budget). This means that the random feature model does not lead to utility loss. Moreover, the excessive risk gaps across different groups are similar for the random feature model, especially when the privacy budget $\epsilon_{p}$ is small, which indicates that better fairness may be achieved by using the private random feature model. In addition, we observe that the excessive risk gap decreases as the privacy budget $\epsilon_{p}$ increases. The same phenomenon has been observed in [28]. This is likely due to the fact that larger $\epsilon_{p}$ values require smaller $\sigma$ values, and thus, as shown in equation (9), the excessive risk gap decreases for a fixed Hessian trace. In summary, we believe that better fairness may be achieved by using the random features model and it is thus superior to feature normalization.
4.3 Statistical Parity
In this section, we adopt the notation of statistical parity as the fairness metric and compare the output distributions of our proposed DP random features model and of the private regularized linear regression model. We use the medical cost and wine quality datasets for numerical experiments as well.
Experiment setup. We keep similar data processing procedures and experiment settings as in Section 4.2. Instead of downsampling and splitting data samples to training and test sets, we use the whole dataset to train the models. Moreover, we fix the privacy budget $\epsilon_{p}=0.5$ for the medical cost example and $\epsilon_{p}=0.05$ for the wine quality example.
Figure 9: Medical cost with sensitive feature ”smoker”: distributions of targets and of the outputs of regressors. From left to right: targets, DP random features model, regularized least-squares, DP regularized least-squared.
Our results are presented in Figure 9 and Table 2. In Figure 9, we depict the distributions of targets in the dataset, outputs of DP random features model, outputs of regularized least-squares method, and DP regularized least-squares method. The result shows that the outputs of the DP random features model conditioned on different sensitive groups have similar distributions. Furthermore, we report estimated statistical parity scores, as defined in Definition 5. To estimate the Kolmogorov-Smirnov distance of output distributions conditioned on different groups, we take a grid of resolution 500 over the domain and estimate the probability that outputs are less than the grid points. Then we take the maximum to get the estimated Kolmogorov-Smirnov distance. The experiments are repeated 20 times over the sampling of random features and the DP mechanism. From Table 2, we observe that the DP random features model can reduce the statistical parity score even if the score is high in the original targets. Notice that our method does not require post-processing to have the same output distributions for different groups. In summary, we believe that better statistical parity may be achieved by using the random features model.
Dataset Targets DP-RF RLS DP-RLS
Medical Cost (“sex”) 0.06 $117\pm 0.047 0.0$74 $0.427\pm 0.233$
Medical Cost (“smoker”) $0.873$ $0.410\pm 0.098$ 0.999 $0.516\pm 0.302$
Wine Quality 0.130 $0.028\pm 0.009$ 0.279 $0.312\pm 0.239$
Table 2: Summary of estimated statistical parity scores.
@@ 5 Discussion
In this paper, we established theoretical guarantees for the differentially private random feature model. The random feature model is a widely used surrogate model for differentially private learning with kernels. Apart from the current literature where people focused on the regularization regime, we focused on the overparametrized regime and studied the min-norm interpolation estimator. Our setting is closer to the modern deep learning setting where the number of trainable parameters is much larger than the number of samples and explicit regularization is not added. Our privacy guarantee follows the standard Gaussian mechanism and the derivation of the generalization error bound relies on tools from random matrix theory. We also show that the proposed DP random features model seems to not suffer disparate impact. Additionally, numerical experiments showed that our method has better generalization performance under the same privacy budget compared with the current differentially private random feature method, and has the potential to reduce disparate impact.
We conclude with some directions for future work. First, it would be interesting to study differentially private shallow neural networks and kernel methods through the random features model. Moreover, empirical privacy evaluation on how random features models leak information about the individual data records deserves future investigation. Finally, we can consider other fairness metrics, such as Equalized Odds. For the fairness results, our focus was on the computational side, and we expect theoretical guarantees on the trade-offs between privacy, fairness, and utility to be interesting future work.
@@
In the appendix, we give a full proof of Theorem 11, which bounds the generalization error of non-private random features model. Lemmas 14 and 15 are used to bound approximation error, which addresses the difference between target function and the ”best” random features model.
@@
In the next lemma, we show the decay rate of the coefficients $\mathbf{c}^{*}$ in the ”best” random features model.
@@
With the above three lemmas at hand, we are ready to present the proof of the generalization error bound of non-private random features model.
@@
We first give a bound of $\|f-f^{*}\|_{L^{2}(\mu)}$. As a consequence of Lemma 14 and 15, we can conclude that
@@
$\|f-f^{*}\|_{L^{2}(\mu)}=\left(\int_{\mathbb{R}^{d}}|f(\mathbf{x})-f^{*}(\mathbf{x})|^{2}d\mu(\mathbf{x})\right)^{1/2}\leq\frac{14\log(2/\delta)}{\sqrt{N}}\|f\|_{\rho}.$ (15)
@@
$v(\mathbf{u}_{1},\dots,\mathbf{u}_{m})$ $=\|f^{*}-f^{\sharp}\|^{2}_{L^{2}(\mu)}-\frac{1}{m}\sum_{j=1}^{m}\left|f^{*}(\mathbf{u}_{j})-f^{\sharp}(\mathbf{u}_{j})\right|^{2}$
@@
$\leq$ $\frac{2N}{m}\|\mathbf{c}^{*}-\mathbf{c}^{\sharp}\|_{2}^{2}=\Delta_{v},$